### v1

# Towards Verifiable Text Generation with Evolving Memory and Self-Reflection

Abstract
Large Language Models (LLMs) face several challenges, including the tendency to produce incorrect outputs, known as hallucination. An effective solution is verifiable text generation, which prompts LLMs to generate content with citations for accuracy verification. However, verifiable text generation is non-trivial due to the focus-shifting phenomenon, the dilemma between the precision and scope in document retrieval, and the intricate reasoning required to discern the relationship between the claim and citations. In this paper, we present VTG, an innovative approach for Verifiable Text Generation with evolving memory and self-reflection. Vtg maintains evolving long short-term memory to retain both valuable documents and up-to-date documents. Active retrieval and diverse query generation are utilized to enhance both the precision and scope of the retrieved documents. Furthermore, Vtg features a two-tier verifier and an evidence finder, enabling rethinking and reflection on the relationship between the claim and citations. We conduct extensive experiments on five datasets across three knowledge-intensive tasks and the results reveal that Vtg significantly outperforms existing baselines.
Towards Verifiable Text Generation with Evolving Memory and Self-Reflection
Hao Sun1, Hengyi Cai2, Bo Wang4, Yingyan Hou2 Xiaochi Wei3, Shuaiqiang Wang3, Yan Zhang1, Dawei Yin3 1Peking University,2Chinese Academy of Sciences 3Baidu Inc, 4Beijing Institute of Technology sunhao@stu.pku.edu.cn
1 Introduction
Large Language Models (LLMs) Scao et al. (2022); Taylor et al. (2022); Chowdhery et al. (2022) have showcased remarkable performance across a spectrum of downstream tasks recently. Despite their advancements, LLMs often generate responses laden with hallucinated facts and inaccurate information Ji et al. (2023); Shuster et al. (2021), undermining their reliability.
To enhance the reliability of LLMs and mitigate the issue of hallucination, it is crucial to encourage these models to provide citations linked to relevant documents in their responses. For example, as shown in Figure 1, a response to the question “Does drinking coffee have health benefits?” is more credible if it cites authentic sources supporting the statement about the reduced risk of type 2 diabetes, compared to a response without any evidence. This approach not only verifies the model’s outputs but also bolsters the authenticity of the generation Gao et al. (2023b).
Figure 1: Given a question, the system generates text while providing citing documents from a large corpus.
However, creating high-quality, well-cited responses poses significant challenges. Firstly, the task of verifiable text generation often involves long text generation, where the focus of the content changes over time, characterized by focus-shifting phenomenon. This dynamic poses challenges in consistently aligning statements with the appropriate evidential references. For instance, as depicted in Figure 1, the discussion may evolve from the health benefits of moderate coffee consumption to the adverse effects of excessive consumption. Such shifts demand a dynamic adaptation in the document pool to appropriately support the shifting focus of the generation. Secondly, striking a balance between the precision and scope of retrieved documents presents a complex challenge for the retrieval system. On the one hand, the task is susceptible to the influence of noisy documents during the citation-sentence alignment process, emphasizing the need for the retrieval system to selectively provide a few highly relevant documents. On the other hand, the intrinsic nature of verifiable text generation calls for a comprehensive collection of documents to enhance credibility. This creates a dilemma: the requirement for precise document retrieval to reduce noise with the need for expansive retrieval to improve the diversity and scope of documents. Lastly, deducing the intricate relationship between a sentence and its potential evidence requires more than just linguistic matching—it demands meticulous consideration. For example, in Figure 1, the third sentence suggests excessive coffee consumption leads to various mental and physical discomforts. The corresponding supporting document, although not explicitly mentioning these discomforts, describes symptoms that are inherently related to them. This situation underscores the importance of in-depth examination to verify that the provided evidence indeed deduces the specific sentence.
When composing text with citations, individuals are capable of gradually gathering the necessary information and adaptively concentrating on the most relevant retrieved references regarding the sentences being written, typically involving deliberate thought and frequent verification. In this paper, we present Vtg, short for Verifiable Text Generation, a framework designed to emulate this human-like process. Vtg operates through iterative generation and verification, utilizing a long short-term memory and a two-tier verifier. Specifically, to address the challenge of focus-shifting in verifiable text generation, Vtg employs an evolving long short-term memory system. This system efficiently stores the most valuable documents in long-term memory and the most recent documents in short-term memory, thereby providing support for the latest sentence. Additionally, Vtg implements an active retrieval process, activated only when the sentence fails to pass the verifier. By doing so, Vtg ensures that the retrieval is necessary, enhancing the precision of the retrieved documents. Before retrieval, LLM is asked to generate diverse queries so that the scope of the retrieved documents can also be improved. Moreover, Vtg integrates a two-tier verifier. In instances where a sentence does not pass the verifier, Vtg activates an evidence finder to retrieve additional supporting evidence. This feature allows the model to rethink and reflect on the relationship between a sentence and its potential evidence, ensuring a thorough and accurate verification process. We conduct experiments on five QA datasets across three knowledge-intensive tasks and the results demonstrate that Vtg obtains superior performance on both citation quality and answering correctness. The effectiveness of the proposed approach is also verified by extensive analytical experiments.
In summary, our main contributions are as follows:
  1. 1.
    We introduce Vtg, a unified framework that guides the generation model using a combination of long short-term memory and a two-tier verifier, providing a self-reflective and nuanced approach for verifiable text generation.
  2. 2.
    Our implementation of long short-term memory effectively captures the most valuable and up-to-date documents, significantly addressing the focus-shifting issue. The active retrieval mechanism and integration of diverse query generation increase both the precision and scope of the retrieved documents. The two-tier verifier and evidence finder enable in-depth analysis of the relationship between generated sentences and potential evidence.
  3. 3.
    We conduct extensive experiments on five datasets across three knowledge-intensive tasks. The empirical results confirm that our method significantly outperforms existing baselines across various metrics.
2 Method
Figure 2: An illustration of our proposed Vtg. Vtg features an evolving memory system, comprising long-term memory $D_{L}$ and short-term memory $D_{S}$. Starting with $D_{L}$ filled with top-$k$ retrieved documents based on the original question and an empty $D_{S}$, the LLM generates a sentence and corresponding citations using both memories. A two-tier verifier then assesses citation adequacy. Successful sentences lead to citation simplification and integration into $D_{L}$ for future use. Unverified sentences trigger the evidence finder to generate new queries, retrieving documents to refresh $D_{S}$ and aid in sentence regeneration.
0:  Question $q$, document pool $\mathbb{D}$, the Generators, the Verifiers, the Citation Simplifier, the Evidence Finder, the maximum trials $T$, the retriever $R$, the number of initially retrieved documents $k$, the number of generated queries $M$, the number of documents retrieved per query $N$
0:  Output with citations $O$
1:  $t\leftarrow 0$
2:  $O\leftarrow\{\}$
3:  $D_{S}\leftarrow\{\}$
4:  $D_{L}\leftarrow R(q,\mathbb{D},k)$
5:  while TRUE do
6:     $s\leftarrow\operatorname{SentenceGenerator}(O,D_{S}\cup D_{L})$
7:     $C\leftarrow\operatorname{CitationGenerator}(s,D_{S}\cup D_{L})$
8:     if  $s$ is NULL  then
9:        break
10:     end if
11:     if  $\operatorname{GenerationVerifier}(s,C)\to$ TRUE  then
12:        $C\leftarrow\operatorname{CitationSimplifier}(s,C)$
13:        $O\leftarrow O\cup\{s,C\}$
14:        $D_{L}\leftarrow D_{L}\cup C$
15:        $t\leftarrow 0$
16:     else if $\operatorname{MemoryVerifier}(s,D_{S}\cup D_{L})\to$ TRUE then
17:        $C\leftarrow\operatorname{CitationSimplifier}(s,D_{S}\cup D_{L})$
18:        $O\leftarrow O\cup\{s,C\}$
19:        $D_{L}\leftarrow D_{L}\cup C$
20:        $t\leftarrow 0$
21:     else if $t>T$ then
22:        $O\leftarrow O\cup\{s,C\}$
23:        $t\leftarrow 0$
24:     else
25:        $D_{S}\leftarrow\operatorname{EvidenceFinder}(s,M,N)$
26:        $t\leftarrow t+1$
27:     end if
28:  end while
29:  return  $O$
Algorithm 1 The pipeline of Vtg
2.1 Overview
In this section, we illustrate the workflow of Vtg in Figure 2, with its procedural specifics outlined in Algorithm 1. Vtg operates with an evolving memory system: the long-term memory $D_{L}$, which oversees the entire generation process, and the short-term memory $D_{S}$, focused on generating the current sentence. Initially, $D_{L}$ is populated with the top-$k$ retrieved documents, while $D_{S}$ begins empty. The LLM generates the first sentence and its corresponding citation using information from both memories. Subsequently, the generation verifier examines whether the cited documents adequately support the sentence. If the documents are insufficient, the memory verifier then evaluates if the combined memories (i.e., $D_{S}\cup D_{L}$) can deduce the sentence. If the sentence is successfully verified by either verifier, the citation set undergoes simplification and is incorporated into $D_{L}$ for future generation tasks. Conversely, if neither verifier confirms the sentence, the LLM generates diverse queries, which are used to retrieve new documents to refresh $D_{S}$. In cases where a regenerated sentence fails to pass verification after $T$ attempts, the model progresses to the next generation stage. This interactive and iterative process continues until the LLM generates an end token, indicating generation completion.
2.2 Generation with Memory
In verifiable text generation, the content’s focus may shift as the generation progresses. Consequently, the pool of supporting documents must be dynamically adjusted to ensure that the documents can deduce the latest sentence. This variability presents a challenge when generating sentences and citations simultaneously due to fluctuating reference indices, which could potentially cause confusion during the model’s generation process. Moreover, insights from previous research Juneja et al. (2023) suggest that tasks requiring distinct skills are more effectively handled through specialized modules instead of a single, multifunctional unit. Based on these insights, Vtg adopts a divided approach: segregating the generation task into distinct sentence generation and citation generation processes.
Sentence Generator
aims to complete unfinished sentences based on the documents from both long-term memory $D_{L}$ and short-term memory $D_{S}$. Initially, long-term memory $D_{L}$ is populated with top-$k$ documents selected based on their relevance to the original question. This ensures that information closely aligned with the original question is permanently retained in memory, providing a consistent guide throughout the entire generation process. In contrast, short-term memory $D_{S}$ starts as an unpopulated set, ready to be dynamically filled as the generation process unfolds.
Citation Generator
aims to source supporting evidence for the current sentence from both long-term memory $D_{L}$ and short-term memory $D_{S}$. This focused approach allows the citation generator to concentrate on validating the current sentence without being distracted by unrelated context.
2.3 Citation Verification and Simplification
Model hallucination in LLM outputs often leads to inaccuracies and distortions Shuster et al. (2021); Ji et al. (2023). One approach to mitigate this issue is by grounding the model’s outputs in specific documents. However, even citations generated by LLM can be subject to hallucinations, potentially leading to unreliable or irrelevant references. To mitigate this issue, Vtg employs a two-tier verification system comprising the generation verifier and the memory verifier. The generation verifier assesses whether the cited documents logically support the sentence, while the memory verifier evaluates the support provided by the entire memory set. Upon successful verification by either component, the citations undergo simplification to remove any unnecessary references.
Generation Verifier
examines whether the citation set logically deduces the sentence. If successful, the citation set advances to the citation simplifier for refinement. However, if the verification fails, two possibilities arise: 1) The sentence is logically consistent with the full memory set, but the generated citation set is inaccurate. In this scenario, the citation simplifier will be used to remove unnecessary citations from the full memory set. 2) The sentence lacks support from the full memory set, suggesting a possible hallucination. Here, the evidence finder is activated to seek relevant information for regenerating the sentence.
Memory Verifier
examines the potential fabrication of sentences by analyzing whether the full memory set logically deduces the current sentences. If the full memory set deduces the current sentences, it points to a potential hallucination in the citation process, triggering the citation simplifier to refine the citation set. On the other hand, if the full memory set fails to support the current sentence, it suggests a possible fabrication by the LLM, necessitating the employment of the evidence finder to search for new evidence.
Citation Simplifier
is designed to align with user preferences for citation elimination, specifically focusing on reducing unnecessary references. It works by iteratively reviewing each citation in the set, temporarily removing one, and assessing if the sentence remains well-supported without it. Redundant citations, that do not contribute to the sentence’s verification, are permanently removed. This iterative process ensures the final citation set is concise and relevant, retaining only essential citations for sentence verification. These citations, after verification and simplification, become part of $D_{L}$ to guide future sentence generations.
2.4 Evidence Finder
When the citation set fails to pass the memory verifier, indicating potential factual inaccuracies in the generated sentence, the evidence finder is utilized to retrieve relevant documents for verification.
Traditional query-generation methods, which typically rely solely on the current sentence, can lead to ambiguity, particularly with sentences containing pronouns or unclear references. To overcome this, Vtg introduces a context-aware query generation approach. This method enhances query generation by integrating the current sentence, the original question, and the preceding unfinished sentence into the prompt. Furthermore, we prompt LLMs to formulate queries that explore various aspects of the current sentence. This comprehensive approach ensures the generated queries are highly related and diverse, thereby expanding the knowledge scope of the retrieved documents.
Once queries are generated, a retriever is employed to collect relevant documents for each query. These documents then refresh the short-term memory $D_{S}$, replacing it with new information. This crucial step ensures the short-term memory remains up-to-date and relevant, thereby enabling Vtg to provide precise citations for the latest sentence.
3 Experiment
Datasets Wikihop WebQ NQ
Correct Citation Correct Citation Correct Citation
Metrics EM F1 Rec Prec F1 EM Rec Prec F1 EM Rec Prec F1
Vicuna-13B
Vanilla 23.40 21.98 29.55 22.25 25.39 55.80 67.66 60.66 63.97 54.80 71.39 61.71 66.20
Summ 23.80 21.64 32.09 26.55 29.06 59.20 60.59 65.60 63.00 56.00 53.57 60.44 56.80
Snippet 21.40 19.53 28.76 25.96 27.29 57.40 53.79 56.62 55.17 57.00 55.84 58.33 57.06
Rerank 22.60 21.13 47.03 47.53 47.28 56.40 89.93 76.33 82.57 56.20 83.56 73.57 78.25
Vtg 25.60 23.27 55.36 49.59 52.32 60.00 92.16 86.51 89.25 58.00 88.69 82.02 85.22
Text-Davinci-003
Vanilla 33.00 33.01 40.46 28.30 33.30 67.50 63.78 58.97 61.28 62.50 60.48 55.56 57.92
Summ 28.00 29.72 19.50 15.98 17.57 67.50 66.40 58.32 62.10 60.00 58.58 52.49 55.37
Snippet 31.00 31.31 23.69 20.01 21.70 65.50 68.35 56.83 62.06 59.00 60.73 53.81 57.06
Rerank 32.67 33.09 56.13 45.22 50.09 67.00 73.72 64.90 69.03 61.50 71.30 63.44 67.14
Vtg 41.50 40.19 63.89 57.65 60.61 68.00 93.00 88.72 90.81 63.00 91.85 86.59 89.14
Table 1: Comparisons between VTG and baselines on Multi-hop QA task and Open-domain QA task.
Datasets ASQA ELI5 Overall
Correct Citation Correct Citation Correct Citation
Metrics EM D-F1 R-L DR Rec Prec F1 Claim Rec Prec F1 EM F1
Vicuna-13B
Vanilla 32.00 27.52 33.53 30.53 72.78 62.09 67.01 12.20 59.79 48.26 53.41 35.64 55.19
Summ 34.44 27.47 35.78 31.63 67.25 63.78 65.47 12.47 59.24 53.36 56.15 37.18 54.09
Snippet 37.62 27.54 36.84 32.19 51.38 53.34 52.34 13.93 40.68 38.19 39.40 37.47 46.25
Rerank 37.14 28.21 32.18 30.20 88.29 75.74 81.53 11.67 73.80 61.12 66.86 36.80 71.30
Vtg 41.92 30.53 37.87 34.20 89.15 82.57 85.73 14.73 81.50 72.16 76.55 40.05 77.81
Text-Davinci-003
Vanilla 40.25 31.47 35.81 33.64 58.13 55.17 56.61 13.43 58.66 47.40 52.43 43.34 52.31
Summ 39.90 31.17 37.26 34.22 59.11 53.77 56.31 13.76 49.09 44.82 46.86 41.83 47.64
Snippet 40.75 30.78 39.07 34.93 62.67 55.34 58.78 12.77 51.71 41.40 45.98 41.80 49.12
Rerank 39.55 29.94 39.38 34.66 75.83 69.81 72.70 14.76 76.21 61.67 68.17 43.10 65.43
Vtg 41.53 31.64 39.45 35.55 86.70 79.95 83.19 16.67 82.63 71.56 76.70 46.14 80.09
Table 2: Comparisons between VTG and baselines on Long-form QA task and overall performance.
3.1 Baselines
For an equitable comparison, we have selected four best-performing baseline methodologies as proposed in ALCE Gao et al. (2023b), which include Vanilla, Summ, Snippet and Rerank. Each of these methods incorporates multiple demonstrations within the initial prompt to facilitate the process of generating responses.
Vanilla.
This configuration involves providing the LLM with the top-$k$ ranked documents. The LLM is then tasked with generating responses that appropriately include citations.
Summ.
In this approach, the LLM is required to synthesize relevant information from the top-$K$ ranked documents. After summarizing these documents, the condensed text is integrated into the prompt. The LLM is then instructed to create texts that incorporate citations, drawing from this summarized content.
Snippet.
In this setup, the LLM is instructed to extract relevant snippets from the top-$K$ ranked documents. These concise documents are subsequently utilized in the prompt, with the aim for the LLM to create text that includes citations, drawing from these brief extracts.
Rerank.
This methodology entails a two-stage process. Initially, the LLM generates four distinct responses using high temperature. Thereafter, each response undergoes an evaluation for citation recall. The response with the highest citation recall score is then chosen as the final output.
3.2 Datasets and Evaluation
We assess the effectiveness of our methods on five datasets across three knowledge-intensive tasks, as detailed in Table 4 of Appendix A. For all datasets, our evaluation criteria encompass both the answer correctness and citation quality of model outputs.
Multihop QA
entails answering complex questions that necessitate multiple retrieval and reasoning steps Yang et al. (2018); Ho et al. (2020). We employ the 2WikiMultihopQA dataset Ho et al. (2020), which consists of 2-hop complex questions derived from Wikipedia, requiring skills in composition, comparison or inference.
In line with Jiang et al. (2023), LLMs are prompted to provide the final answer, which is then evaluated against the reference answer using answer-level exact match (EM), and token-level F1, precision, and recall metrics.
Long-form QA
aims to generate detailed answers to complex questions Fan et al. (2019a, b), we utilize the ASQA Stelmakh et al. (2022) and ELI5 Fan et al. (2019b) datasets. ASQA focuses on ambiguous questions requiring comprehensive answers covering multiple interpretations. ELI5, on the other hand, deals with complex questions demanding lengthy, in-depth answers backed by multiple documents.
For ASQA, we apply the metrics outlined in Stelmakh et al. (2022), including EM, a soft match using a RoBERTa-based QA model (Disambig-F1), ROUGE Lin (2004), and a combined DR score. In the case of ELI5, we adhere to the evaluation criteria of Gao et al. (2023b), focusing on whether the model’s predictions address the sub-claims of the gold-standard answer.
Open-domain QA
requires leveraging external knowledge for answering questions. Our datasets of choice are NQ Kwiatkowski et al. (2019) and WebQ Berant et al. (2013).
Following the methodology of Sun et al. (2023), LLMs are prompted to generate the final answer, which is then compared with the reference answer using answer-level exact match (EM), and token-level F1.
Verifiability Evaluation
To evaluate the citation quality of responses, we employ the approach of Gao et al. (2023b), focusing on calculating Citation Recall, Citation Precision, and the combined Citation F1 score. Citation Recall examines whether the output is fully supported by the cited documents, while Citation Precision assesses the redundancy of the citations included.
3.3 Implementation Details
To ascertain the generalizability of our method, we conducted experiments using different sizes of Large Language Models (LLMs). Specifically, we utilized Vicuna-13B-v1.5-16k Zheng et al. (2023) [^1] https://huggingface.co/lmsys/vicuna-13b-v1.5-16k and Text-Davinci-003 Ouyang et al. (2022) [^2] https://api.openai.com/v1/completions as of October 2023 for evaluation, respectively.
In alignment with prior research Trivedi et al. (2022); Jiang et al. (2023), we randomly subsampled 500 examples for Vicuna-13B-v1.5-16k and 200 examples for Text-Davinci-003 from each dataset due to the cost of running experiments. For the evaluation of verifiability and inference tasks in both the Rerank and Vtg models, we employed the TRUE [^3] https://huggingface.co/google/t5_xxl_true_nli_mixture model (Raffel et al., 2020), a T5-11B model fine-tuned on a collection of NLI datasets, to automatically examine whether the cited documents entail the model generation.
3.4 Main Results
In this section, we delineate our primary findings derived from extensive analyses conducted on five diverse datasets, as outlined in Tables 1 and 2. These investigations yield several key observations:
First, our proposed Vtg method consistently outperforms other approaches across various datasets and metrics when applied to LLMs of different sizes. Notably, Vtg enhances citation quality significantly, achieving a remarkable 20% and 9% improvement over our strongest competitor Rerank, using Text-Davinci-003 and Vicuna-13B-v1.5-16k, respectively. Furthermore, Vtg’s capacity for verifiable generation substantially enhances correctness, evidenced by an approximate 6% overall improvement compared to the leading baseline across both LLMs.
Second, among the evaluated baselines, Rerank stands out in citation quality, primarily due to its multiple sampling strategy that increases the likelihood of high-quality outputs. However, its performance in terms of correctness is inconsistent across different datasets. For instance, Rerank’s effectiveness is notably lower in ASQA’s EM using Text-Davinci-003 and ELI5’s Claim using vicuna-13B-v1.5-16k. In contrast, Vtg demonstrates consistent improvement across all metrics and datasets, highlighting its robustness and reliability.
Third, the comparison between two different LLMs, i.e., vicuna-13B-v1.5-16k and Text-Davinci-003, reveals intriguing distinctions. Methods like Summ and Snippet, which integrate a broader range of documents, achieve greater correctness with vicuna-13B-v1.5-16k but falter with Text-Davinci-003. This disparity can be attributed to Text-Davinci-003’s comparatively extensive knowledge base, as evidenced by the significant improvement in correctness when applying identical methodologies to both LLMs. Consequently, Summ and Snippet may introduce extraneous information in the context of Text-Davinci-003. Moreover, these methods underperform in citation quality with both LLMs, potentially due to the fact that their reliance on simplified document versions challenges the verification of document support for the claims.
Datasets Correct Citation
EM F1 Rec Pre F1
Vtg 25.60 23.27 55.36 49.59 52.32
-w/o QG 21.20 18.98 48.97 41.54 44.95
-w/o Verifier 21.00 17.72 33.09 20.07 24.98
-w/o Simplifier 24.40 22.23 46.38 29.47 36.04
Table 3: Ablation Study on Wikihop.
Figure 3: The performance change over different hyper-parameters on ASQA.
3.5 Ablation Study
In this section, we evaluate the impact of each component in our model using the Wikihop dataset. By removing components one by one, we observe their individual contributions to performance. “w/o QG” replaces query generation with direct retrieval using the original sentence. “w/o Verifier” excludes the two-tier verifier during generation, and “w/o Simplifier” removes the citation simplification.
Results in Table 3 show that removing any component decreases performance, highlighting their importance. Notably, the removal of the verifier leads to the most significant drop, as it results in the generation of sentences without external validation, potentially introducing hallucinations into the generation. On the other hand, omitting the simplifier has a less pronounced effect on correctness, mainly affecting citation quality through reduced precision. Removing QG limits the model’s ability to retrieve a broader range of relevant documents, leading to performance degradation.
3.6 Hyper-parameter Analysis
In this section, we examine the impact of various hyper-parameters on Vtg’s performance on ASQA dataset, with a focus on the number of documents retrieved per query $N$, the number of generated queries $M$, and the maximum trials $T$, as illustrated in Figure 3. We adjust these parameters one at a time to observe their individual effects. Our findings are as follows:
For the number of documents retrieved per query $N$, initially increasing $N$ enhances citation quality but reduces answer correctness. This trend suggests that a larger pool of documents provides more citation options for the LLM, but also introduces potentially noisy information, impacting model generation. When $N$ becomes very large, there’s a consistent decrease in both citation quality and answer correctness, likely due to the overwhelming amount of information confusing the model.
For the number of generated queries $M$, the optimal performance for both answer correctness and citation quality occurs at $M$=2. Beyond this, performance declines. When $M$ is small, increasing it broadens the knowledge scope for document retrieval, thereby improving outcomes. However, a larger $M$ can lead to less relevant queries, introducing noise and thus degrading model performance.
For the maximum trials $T$, a higher $T$ correlates with improved performance, as it allows the model more attempts to generate a sentence that passes verification. This approach provides the model with more chances for a successful generation. However, an increase in $T$ also results in higher token consumption, indicating a need to balance effectiveness with computational cost. Adjusting $T$ can help find an equilibrium between model performance and computational cost.
Figure 4: Token Consumption Comparison on NQ.
3.7 Token Consumption Analysis
This section presents a comparative analysis of token consumption by different methods on the NQ dataset, as detailed in Figure 4.
As we can find, Summ, Snippet, and Vanilla incur nearly identical token costs per query. Rerank, which samples responses four times, incurs the highest token cost, albeit yielding superior citation quality compared to other baselines. Contrarily, our method Vtg demonstrates a lower token cost than Rerank while significantly enhancing citation quality, thereby underscoring the efficiency of our approach. Furthermore, as discussed in Section 3.6, the token cost in our method can be modulated through the max trials hyperparameter, providing a versatile option for balancing performance with resource expenditure.
4 Related Work
4.1 Retrieval-augmented LLMs.
Retrieval-augmented LLMs aim to provide extra documents to the LLM, which has been proven useful in many knowledge-intensive tasks. Among the existing studies, Shi et al. (2023); Wang et al. (2023b); Zhang et al. (2023); Yu et al. (2023a, c) propose to retrieve only once at the beginning. Other works Qian et al. (2023); Yu et al. (2023b) propose to retrieve multiple times during generation, which offers flexibility of when and what to search. For example, Jiang et al. (2023) propose to retrieve when the generated sentence contains low confidkkence tokens. Ram et al. (2023) propose to refresh the retrieved document every n token, which is demonstrated to be more effective than retrieving only once. Wang et al. (2023a); Asai et al. (2023); Zhao et al. (2023) propose to retrieve only when the LLMs need to.
Among the existing studies, retrieve on-the-fly methods are the closest to ours. However, these methods do not provide referenced documents for the generated sentences, potentially reducing the reliability of the generated content.
4.2 Verifiable Text Generation
Verifiable Text Generation aims to generate content with supporting documents, which has been attracting attention in recent years. For example, Liu et al. (2023b); Qin et al. (2023); Nakano et al. focus on training LLMs to browse web pages and answer questions with evidence. Gao et al. (2023a) introduced the research-and-revision (RARR) method for retrieving evidence for LLM outputs. Li et al. (2023) incorporated knowledge graphs as an evidence source. Other works mainly focus on evaluation Liu et al. (2023a). For example, Rashkin et al. (2023) propose Attributable to Identified Sources (AIS) for human evaluation. Gao et al. (2022) define auto-AIS to approximate human AIS judgments. Gao et al. (2023b) propose ALCE to automate the evaluation of the citation quality. Min et al. (2023) proposed FactScore for evaluating the verifiability of generated facts.
However, these methods do not effectively address the focus-shifting problem. In this paper, we propose incorporating a long short-term memory to retain important documents in the long-term memory and current documents in the short-term memory.
5 Conclusion
In this paper, we introduce Vtg, a novel framework tailored to address the complexities of verifiable text generation. Central to its design is an evolving long short-term memory, which adaptively keeps both valuable documents and up-to-date documents. Through the integration of active retrieval mechanisms and diverse query generation, Vtg skillfully enhances both the precision and scope of the document retrieval process. The two-tier verifier coupled with an evidence finder facilitates a deeper analysis and reflection on the relationship between claims and citations. Extensive experiments on five datasets across three knowledge-intensive tasks verify the effectiveness of our method.
Appendix A Datasets and Settings
Datasets and experimental settings are summarized in Table 4.
Settings 2WikiMultihopQA ASQA ELI5 NQ WebQ
Ho et al. (2020) Stelmakh et al. (2022) Fan et al. (2019b) Kwiatkowski et al. (2019) Berant et al. (2013)
Dataset statistics
Task Multihop QA Long-form QA Long-form QA Open-domain QA Open-domain QA
#Examples (Vicuna) 500 500 500 500 500
#Examples (Davinci) 200 200 200 200 200
Evaluation settings
Correctness Metrics EM, Token-level F1, Pre, and Rec EM, Disambig-F1, ROUGE, DR Claim EM, F1 EM, F1
Citation Metrics Prec, Rec, F1
Retrieval settings
Corpus Wikipedia Wikipedia Wikipedia Wikipedia Wikipedia
Retriever DPR DPR DPR DPR DPR
Table 4: Statistics and experimental settings of different tasks/datasets.
Appendix B Prompts
Instruction: Write an accurate, engaging, and concise answer for the given question using only the provided search results (some of which might be irrelevant). Question: {Question} Document: {Document} Answer:
Instructions: You will be provided with a sentence and several related documents. Your task is to directly append citation annotations to the sentence using these documents without changing the sentence. When citing documents, use [1][2][3]. Cite at least one document and at most three documents. If multiple documents support the sentence, only cite a minimum sufficient subset of the documents. Document: {Document} Sentence: {Sentence} Sentence with citation:
Given the claim: claim, please generate up to num questions that can help verify the claim with the following constraints: 1. You should output no more than three questions. 2. The generated questions should be diverse and focus on different aspects of the given question. Generated questions:
Prompt B.4: Summ Instruction
Summarize the following document within 50 words with the question of interest {Question} Return "irrelevant" if the document is “irrelevant" to the question. Try to keep all the important dates, numbers, and names. Title: {Title} Text: {Text} Summary:
Prompt B.5: Snippet Instruction
Given the following passage and the question {Question}, extract a useful span from the passage that can answer the question. Resolve all the coreference issues to make the extracted span understandable and standalone. If the passage is not helpful for answering the question, return “irrelevant". If there are multiple spans, merge them and only output one paragraph. Title: {Title} Text: {Text} Extracted span: